# Learning Mean-Field Games through Mean-Field Actor-Critic Flow

### v1

@@ Appendix C Proofs for the critic
$\leq C\left(\beta_{\mu}^{2}\,d_{\beta}(\mu^{\tau},\rho^{\mu^{\tau},\alpha^{\tau}})^{2}+\beta_{a}^{2}\left\|\nabla_{\alpha}H(t,x,\mu_{t},\alpha^{\tau}(t,x),-\mathcal{G}^{\tau}(t,x))\right\|_{\mu^{\tau},\alpha^{\tau}}^{2}\right),$
where the equality follows from Lemma A.15 and (3.11a). Substituting (D.12) into (D.11) gives
@@
Here, we record this constant $C_{T}$ to determine the weight $\lambda_{T}$. Substituting (D.10) and (D.13) into (D.9), we get
@@
Let $\lambda_{T}=\frac{1}{4C_{T}}$. Combining (D.8) and (D.14) yeilds the evoluation of the Lyapunov functional
@@
where $c_{\mu}=\frac{1}{2}$.
@@ Appendix E Baseline derivations of models in Section 6
In this appendix, we derive the mean-field equilibria for the systemic risk model and the optimal execution problem, serving as analytical baselines for the numerical comparisons presented in Section 6.
@@
with terminal condition $v(T,x)=\frac{c}{2}(x-m_{T})^{2}$. Using a quadratic ansatz $v(t,x)=\tfrac{1}{2}\eta_{t}(x-m_{t})^{2}+\xi_{t}$, where $\eta,\xi$ are deterministic measurable functions of time. Mimizing over $\alpha$ yields the optimal control
@@
To evaluate the Lyapunov function of the actor (4.3), we need analytic expressions for the control $\alpha^{\mu,*}$, which requires calculations of $V^{\mu,*}$ for any fixed flow of measure $(\mu_{t})$. Given $(m_{t}=\int x\mu_{t}(\mathrm{d}x))$, the function $V^{\mu,*}$ satisfies the HJB equation (E.1). Using a quadratic ansatz $V^{\mu,*}(t,x)=\tfrac{1}{2}\eta^{\mu}_{t}x^{2}+\rho^{\mu}_{t}x+\xi^{\mu}_{t}$, where $\eta^{\mu},\rho^{\mu},\xi^{\mu}$ are deterministic measurable functions of time, we get
@@
Let $m_{t}$ again be the mean of $\mu_{t}$ for any $t\in[0,T]$ and $v:=V^{\mu,*}$ the optimal value function of the representative agent under the given flow of measure $(\mu_{t})$. Since the optimal execution problem is an extended MFG, $\mu_{t}$ denotes a measure on the action space, while the state population distribution is denoted by $\nu_{t}$ with mean $p_{t}=\int x\nu_{t}(\mathrm{d}x)$.
@@ Appendix F Additional numerical experiments for MFAC
In this appendix, we present additional numerical results for the MFAC algorithm applied to the flocking model, complementing the discussion in Section 6.3. Unless otherwise stated, all model parameters are identical to those in Section 6.3, and all hyperparameters follows Appendix G. The only modification concerns the value of the parameter $\beta$.
@@ Appendix G Hyperparamters for numerical experiments
This appendix summarizes the hyperparameters used to produce the numerical results in Section 6 and Appendix F.

### v2

@@ Appendix C Proofs for the critic
$\leq C\left(\beta_{\mu}^{2}\,d_{\beta}(\mu^{\tau},\rho^{\mu^{\tau},\alpha^{\tau}})^{2}+\beta_{a}^{2}\left\|\nabla_{\alpha}H(t,x,\mu_{t},\alpha^{\tau}(t,x),-\mathcal{G}^{\tau}(t,x))\right\|_{\mu^{\tau},\alpha^{\tau}}^{2}\right).$
Substituting (D.12) into (D.11) provides
@@
Here, we record $C_{T}$ for the specification of $\lambda_{T}$. Substituting (D.10) and (D.13) into (D.9) yields
@@
Since $\lambda_{T}\leq\frac{1}{4C_{T}}$, combining (D.8) and (D.14) yields
@@
where $c_{\mu}=\frac{1}{2}$. This concludes the proof. ∎
@@ Appendix E Baseline derivations of models in Section 6
In this section, we derive the mean-field equilibria for the systemic risk model and the optimal execution problem, serving as analytical baselines for the numerical comparisons presented in Section 6.
@@
with terminal condition $v(T,x)=\frac{c}{2}(x-m_{T})^{2}$. We adopt a quadratic ansatz $v(t,x)=\tfrac{1}{2}\eta_{t}(x-m_{t})^{2}+\xi_{t}$, where $\eta,\xi$ are deterministic measurable functions of time. Minimizing over $\alpha$ yields the optimal control
@@
To evaluate the Lyapunov function of the actor (4.3), we need analytic expressions for the control $\alpha^{\mu,*}$, which requires calculations of $V^{\mu,*}$ for any fixed flow of measure $(\mu_{t})$. Given $m_{t}=\int x\,\mathrm{d}\mu_{t}(x)$, the function $V^{\mu,*}$ satisfies the HJB equation (E.1). Using a quadratic ansatz $V^{\mu,*}(t,x)=\tfrac{1}{2}\eta^{\mu}_{t}x^{2}+\rho^{\mu}_{t}x+\xi^{\mu}_{t}$, where $\eta^{\mu},\rho^{\mu},\xi^{\mu}$ are deterministic measurable functions of time, we get
@@
Let $m_{t}$ be the mean of $\mu_{t}$ for any $t\in[0,T]$ and $v:=V^{\mu,*}$ be the optimal value function of the representative agent under the given flow of measure $(\mu_{t})$. Since the optimal execution problem is an extended MFG, $\mu_{t}$ denotes a measure on the action space, while the state population distribution is denoted by $\nu_{t}$ with mean $p_{t}:=\int x\,\mathrm{d}\nu_{t}(x)$.
@@ Appendix F Additional numerical experiments for MFAC
In this appendix, we present additional numerical results for the MFAC algorithm applied to the flocking model, complementing the discussion in Section 6.3. Unless otherwise stated, all model parameters are identical to those in Section 6.3, and all the hyperparameters follow Appendix G. The only modification concerns the value of the parameter $\beta$.
@@ Appendix G Hyperparamters for numerical experiments
This section summarizes the hyperparameters used to produce the numerical results in Section 6 and Appendix F.